# Invariant derivations and trace bounds

### v3

@@ 1. Introduction
for all $g\in G$, where $\chi(1)=\dim V$ and $\operatorname{Re}\colon\mathbb{C}\to\mathbb{R}$ gives the real part of a complex number. As $g$ varies over $G$, the maximum $\chi(1)$ of $\operatorname{Re}\chi(g)$ from (1.1) is trivially achieved, but the minimum need not be, as Serre’s bounds show. For example, for $G=\operatorname{SU}(n)$ and $\rho=\operatorname{Ad}$, the adjoint representation, $\min_{g\in G}\chi(g)$ is asymptotically equivalent to $-\chi(1)/n^{2}$, see Example 3.2. In §17 we observe that $\min_{g\in G}\operatorname{Re}\chi(g)$ is an algebraic number. In §18 we provide some bounds on $\min_{g\in G}\operatorname{Re}\chi(g)$ for all $\chi$.
Finally, Appendix A uses Serre’s trace bounds for compact groups to bound Brauer characters of representations of finite groups over fields of prime characteristic.
@@
We also consider algebraic groups as in Milne’s book [Mil17]. An algebraic group over a field $F$ is a group object in the category of schemes of finite type over $F$. Two examples are $\operatorname{GL}_{n,F}$, the group of $n$-by-$n$ matrices over $F$, and $U(n)_{F}$, the group of $n$-by-$n$ unitary matrices with entries in $F(\sqrt{-1})$. When $F=\mathbb{R}$, these are also real Lie groups. In this paper, we only consider algebraic groups that are affine, and for simplicity we write simply “$G$ is an algebraic group” to mean that $G$ is an affine algebraic group in the sense of Milne’s book. Furthermore, apart from Remark 5.13 and the appendix, we will only consider the case where the field $F$ has characteristic zero, so $G$ will be linear (meaning that it is isomorphic to a Zariski-closed $F$-subscheme of $\operatorname{GL}_{n,F}$ in a way that is compatible with the group operations on $G$ and $\operatorname{GL}_{n,F}$) and smooth, see [Mil17, Prop. 1.41, Th. 3.23, Cor. 4.10]. Because of the assumption of characteristic zero and that $G$ is affine, it is the same to consider these groups from the perspective of [Bor91] or [Spr98]. For any commutative $F$-algebra $R$, we write $G(R)$ for the set $\operatorname{Hom}(\operatorname{Spec}R,G)$ of $R$-points of $G$. The algebraic group $\operatorname{GL}_{1,F}$, i.e., the 1-dimensional torus with $F$-points $F^{\times}$, is typically denoted $\mathbb{G}_{\mathrm{m}}$. We sometimes write $\operatorname{Lie}(G)$ or $\mathfrak{g}$ for the Lie algebra of $G$, and similarly for other algebraic groups.
We say that an algebraic group $G$ is reductive if its unipotent radical is trivial, i.e., if the identity component $G^{\circ}$ is reductive in the sense of [Mil17, §19c]. We say it is simple ([Mil17, Def. 17.97] says “almost simple”) if it is semisimple and non-commutative, and every proper normal algebraic subgroup is finite.
A compact (real) Lie group $G$ can also be viewed as a reductive real algebraic group. The connected components of $G$ are the same, regardless of which viewpoint one takes. (The fact that the two notions coincide on the identity component is [Bor91, 24.6(ii)] and the more general statement follows, see [Con10].) Of course, for non-compact group Lie groups this is typically false.
@@ 2. Serre’s bounds on the trace
Here $\operatorname{Ad}\colon K\to\operatorname{GL}(\operatorname{Lie}(K))$ is the adjoint representation, the action of $K$ on its own Lie algebra by conjugation. This result is proved in §14 as a consequence of the following more detailed result, which considers both the minimums and maximums of $\operatorname{Tr}\operatorname{Ad}$ at the price of restricting to simple compact Lie groups. See [FG98] and [Fre01] for calculations of the value of $\operatorname{Tr}\operatorname{Ad}(x)$ for various $x$, in the case where $K$ is simple of type $E_{8}$.
@@ 2 $-(\operatorname{rank}G)$ 26
In addition to the results we prove here, Serre has recently publicized in [Ser23] his proofs of the other results announced in his 2004 note, so now proofs of all of the announced results are available.
@@ 3. Preliminary observations
Let $G$ be a simple compact Lie group with maximal torus $T$. Suppose $\phi\in\operatorname{Aut}(G)(\mathbb{C})$ normalizes $T$ and acts as $\phi(t)=t^{-1}$ for $t\in T$. Then $\phi$ interchanges the positive and negative roots, so the only diagonal entries of $\operatorname{Ad}(\phi)$ are its action on the torus and we have $\operatorname{Tr}\operatorname{Ad}(\phi)=-(\operatorname{rank}G)$.
@@
The element $g$ is contained in a maximal torus $T$ of $G$. Put $W$ for the Weyl group of $G$ with respect to $T$. Write $f$ as a composition $T\to T/W\to\mathbb{C}$. Suppose $g$ is regular. Then the Jacobian of the first map is invertible at $g$ by Steinberg’s Theorem as in [Ste60] or [Ste65, 6.1, 6.3]. Since $g$ is a critical point of $f$, it follows that the gradient of the map $T/W\to\mathbb{C}$ must vanish at $g$, i.e., ii. ∎
We also need the following small elaboration of results of Steinberg. We provide a proof because we do not know a reference for it in the literature. Every algebraic group $G$ acts on itself by conjugation, and we write $\operatorname{Class}(G)$ for the quotient of $G$ by this action (in order to avoid the notation $G/G$). As a scheme, for $G$ an algebraic group over $F$, $\operatorname{Class}(G)$ is $\operatorname{Spec}F[G]^{G}$.
@@
    If the torus $Z(G)^{\circ}$ is split and the derived subgroup of $G$ is simply connected, then
@@
To verify that the natural map in a is an isomorphism, it suffices to do so when $F$ is algebraically closed.
Let $T^{\prime}$ be a maximal torus in the derived subgroup $G^{\prime}$ of $G$. The identity component $Z(G)^{\circ}$ of the center of $G$ is a torus and $T:=T^{\prime}\cdot Z(G)^{\circ}$ is a maximal torus of $G$. By conjugacy of maximal tori in $G$, we may assume that the given torus $T$ is this $T$. Set $K:=T^{\prime}\cap Z(G)^{\circ}$, $T^{+}:=T^{\prime}\times Z(G)^{\circ}$, and $G^{+}:=G^{\prime}\times Z(G)^{\circ}$. There is a natural diagram
@@
For the left vertical arrow, $T^{+}/W=T/W\times Z(G)^{\circ}$ and $\operatorname{Class}(G^{+})=\operatorname{Class}{G^{\prime}}\times Z(G)^{\circ}$, and these are isomorphic because $G^{\prime}$ is semisimple [Ste65, 6.4]. That is, the left vertical arrow is an isomorphism. The top and bottom arrows are a quotient map by $K$, on which $W$ and $G^{+}$ both act trivially, so the right arrow is also an isomorphism, proving a.
@@
It thus remains to consider the action of $K:=G^{\prime}\cap Z(G)^{\circ}$. Since $Z(G)^{\circ}$ is split, $\operatorname{Hom}(K,\mathbb{G}_{\mathrm{m}})$ is a finite abelian group with trivial action by the Galois group, and thus the action of $K$ induces an $F$-linear grading of $V_{G^{\prime}}$ by $\operatorname{Hom}(K,\mathbb{G}_{\mathrm{m}})$. In addition, the map $\operatorname{Hom}(Z(G)^{\circ},\mathbb{G}_{\mathrm{m}})\to\operatorname{Hom}(K,\mathbb{G}_{\mathrm{m}})$ is surjective, and thus we can rescale each homogeneous component by a character of $Z(G)^{\circ}$ to get a $K$-invariant affine factor. We thus get an isomorphism
$\operatorname{Class}(G)\cong(\mathbb{A}^{r^{\prime}}\times Z(G)^{\circ})/K\cong\mathbb{A}^{r^{\prime}}\times(Z(G)^{\circ}/K)\cong\mathbb{A}^{r^{\prime}}\times G^{\text{ab}}$
as required.
When $G^{\prime}$ is also split, $V_{G^{\prime}}$ is the span of fundamental characters, all of which are homogeneous for the grading by $\operatorname{Hom}(K,\mathbb{G}_{\mathrm{m}})$, and thus the procedure of the previous paragraph multiplies the fundamental characters by characters of $Z(G)^{\circ}$ to obtain new characters that descend to $G$. We thus get an isomorphism of the form
@@
The group $G=\operatorname{U}(2)$ has simply connected derived subgroup $G^{\prime}=\operatorname{SU}(2)$, with $\operatorname{Class}(G^{\prime})\cong\mathbb{A}^{1}$. However, $Z(G)^{\circ}$ is a rank $1$ torus that is not split and we claim that $\operatorname{Class}(G)$ is not isomorphic to the product of an affine space and a torus.
@@ 4. Invariant derivations and critical points
The “only if” direction is trivial. So suppose $t$ is a critical point for $f|_{T}$. The identity component $G^{\circ}$ of $G$ is a union of conjugates of $T$, and thus in particular the conjugates of $\mathfrak{t}$ span $\mathfrak{g}$. For $x\in\mathfrak{g}$, there exists a $g\in G$ and $y\in\mathfrak{t}$ such that $(\operatorname{Ad}g)y=x$ and we have
$xf(t)=((\operatorname{Ad}g)y)f(t)=y(g^{-1}f)(t)=yf(gtg^{-1}).$
By hypothesis, the last term equals $yf(t)$, which is zero because $t$ is a critical point for $f|_{T}$. ∎
@@
$x$ is a critical point for $f$ if and only if $(Df)(x)=0$ for every derivation $D$ (not necessarily $H$-invariant), so the “only if” direction is trivial and we assume $(Df)(x)=0$ for every $H$-invariant $D$. Let $D_{0}$ be any derivation. Since $H$ is finite, there is a $\pi\in F[T]$ such that $\pi(t)=1$ and $\pi$ vanishes on the other elements of the $H$-orbit of $x$. Put $H_{x}$ for the centralizer of $x$ in $H$ and define
@@ 5. Derivations on $T/W$ versus $W$-invariant derivations on $T$
In this section, we work in the context of a connected reductive algebraic group over $\mathbb{C}$ with a maximal torus $T$ and Weyl group $W:=N_{G}(T)/T$. We relate the $W$-invariant derivations on $T$ to derivations on $T/W$. Any $W$-invariant derivation on $T$ takes $W$-invariant functions to $W$-invariant functions, so induces a derivation on $T/W$, giving us a natural map $\operatorname{Der}(\mathbb{C}[T])^{W}\to\operatorname{Der}(\mathbb{C}[T/W])$. Since $\mathbb{C}[T/W]\subset\mathbb{C}(T)$ is étale, a derivation on $T$ is determined by its values on $\mathbb{C}[T/W]$, so this map is injective. This lets us identify $\operatorname{Der}(\mathbb{C}[T])^{W}$ with a submodule of $\operatorname{Der}(\mathbb{C}[T/W])$.
@@
$\operatorname{Der}(\mathbb{C}[T/W])/\operatorname{Der}(\mathbb{C}[T])^{W}\cong\mathbb{C}[t]/(t^{2}-4)$
and in particular that $\operatorname{Der}(\mathbb{C}[T])^{W}$ is a proper submodule of $\operatorname{Der}(\mathbb{C}[T/W])$.
It turns out that when $T/W$ is sufficiently nice (e.g., when the derived subgroup $G^{\prime}$ is simply connected), not only can we show that the relevant submodule of $\operatorname{Der}(\mathbb{C}[T/W])$ is a free submodule, but the proof gives us a reasonably natural way to construct a basis. (A similar construction appears in §B.2 of [OT13], although that material concerns the linear action of $W$, the quotient of a euclidean space by the reflection representation of $W$.)
@@
We take a moment to review some background on differential operators on an affine variety $T$ over $\mathbb{C}$.
@@
A differential operator of order 0 is $\mu_{f}\colon\mathbb{C}[T]\to\mathbb{C}[T]$ given by multiplication by an element $f\in\mathbb{C}[T]$. A differential operator $\mathcal{D}$ of order $k+1$ is a $\mathbb{C}$-linear map $\mathbb{C}[T]\to\mathbb{C}[T]$ such that, for each $g\in\mathbb{C}[T]$, the map $f\mapsto\mathcal{D}(gf)-g\mathcal{D}(f)$ is a differential operator of order $k$.
The collection of differential operators of order $k$ is denoted $\operatorname{Diff}^{2}_{\mathbb{C}[T]/\mathbb{C}}(\mathbb{C}[T],\mathbb{C}[T])$ in [Sta18, Tag 09CH], and it is a $\mathbb{C}$-module. A differential operator of order $k$ is a differential operator of order $K$ for all $K\geq k$. A differential operator of order 1 is a derivation if and only if it annihilates 1 [Sta18, Tag 09CM].
@@
Let $\mathcal{D}$ be a second-order differential operator on an affine variety $T$ over $\mathbb{C}$. Then for $f,g\in\mathbb{C}[T]$:
@@
    $[\mu_{f},\mathcal{D}]+\mu_{\mathcal{D}(f)}-\mu_{\mathcal{D}(1)f}$ is a derivation on $T$.
@@
    The pairing $(f,g)\mapsto\mathcal{D}(fg)-f\mathcal{D}(g)-\mathcal{D}(f)g+fg\mathcal{D}(1)$ is symmetric and is a derivation on $T$ in each slot. It is identically zero if $\mathcal{D}$ is a first-order differential operator.
@@
For b, the pairing is evidently symmetric and $\mathbb{C}$-linear. If $\mathcal{D}$ is first order, then there is an $f^{\prime}\in\mathbb{C}[T]$ such that $\mathcal{D}(fg)-f\mathcal{D}(g)=f^{\prime}g$ for all $g$. Then $\mathcal{D}(f)g-fg\mathcal{D}(1)=f^{\prime}g$ as well and the pairing is zero. Finally, the the image of $(f,g)$ is $-1$ times the derivation in a, so it is a derivation in $g$. ∎
@@
Let $A$ be a nonsingular $W$-invariant quadratic form on $\mathfrak{t}^{*}$. We may view this as an element of $(S^{2}\mathfrak{t})^{W}$, i.e., a Weyl-invariant element of the symmetric square. Since elements of $\mathfrak{t}$ are naturally translation-invariant derivations, $A$ corresponds naturally to a second-order differential operator $\mathcal{D}_{A}$ on $\mathbb{C}[T]$, invariant under $W$ and under translation by $T$.
@@
$\Omega_{\mathbb{C}[T/W]}\to\operatorname{Der}(\mathbb{C}[T/W])$
induced by $M_{A}$ is $\operatorname{Der}(\mathbb{C}[T])^{W}$.
@@
$\Omega_{\mathbb{C}[T]}\cong\operatorname{Der}(\mathbb{C}[T]),$
@@
$\Omega_{\mathbb{C}[T]}^{W}\cong\operatorname{Der}(\mathbb{C}[T])^{W}.$
@@
$\Omega_{\mathbb{C}[T/W]}\to\operatorname{Der}(\mathbb{C}[T/W])$
@@
$\Omega_{\mathbb{C}[T/W]}\to\Omega_{\mathbb{C}[T]}^{W}\cong\operatorname{Der}(\mathbb{C}[T])^{W}\hookrightarrow\operatorname{Der}_{\mathbb{C}[T/W]}$
in which the maps on either side are natural. Moreover, by [Kni73, Cor. to Thm. II.3], the first natural map is an isomorphism, and thus the image of the map induced by $M_{A}$ is $\operatorname{Der}(\mathbb{C}[T])^{W}$ as required. ∎
@@
If the derived subgroup $G^{\prime}$ of $G$ is simply connected, then $\operatorname{Der}(\mathbb{C}[T])^{W}$ is a free module.

### v4

@@ 1. Introduction
for all $g\in G$, where $\chi(1)=\dim V$ and $\operatorname{Re}\colon\mathbb{C}\to\mathbb{R}$ gives the real part of a complex number. As $g$ varies over $G$, the maximum $\chi(1)$ of $\operatorname{Re}\chi(g)$ from (1.1) is trivially achieved, but the minimum need not be, as Serre’s bounds show. For example, for $G=\operatorname{SU}(n)$ and $\rho=\operatorname{Ad}$, the adjoint representation, $\min_{g\in G}\chi(g)$ is asymptotically equivalent to $-\chi(1)/n^{2}$, see Example 3.2. In §17 we observe that $\min_{g\in G}\operatorname{Re}\chi(g)$ is an algebraic number. In §18 we provide some bounds on $\min_{g\in G}\operatorname{Re}\chi(g)$ for all $\chi$ and fixed $G$.
Finally, Appendix A uses results from the main body of the paper to bound Brauer characters of representations of finite groups over fields of prime characteristic.
@@
In most of this paper, we consider objects over a base field $F$ that is either $\mathbb{R}$ or $\mathbb{C}$. In §17 we consider other fields of characteristic zero, and in Remark 5.13 and the appendix we will consider fields of prime characteristic.
@@
We also consider algebraic groups as in Milne’s book [Mil17]. An algebraic group over a field $F$ is a group object in the category of schemes of finite type over $F$. Two examples are $\operatorname{GL}_{n,F}$, the group of $n$-by-$n$ matrices over $F$, and $U(n)_{F}$, the group of $n$-by-$n$ unitary matrices with entries in $F(\sqrt{-1})$. When $F=\mathbb{R}$, these are also real Lie groups. In this paper, we only consider algebraic groups that are affine, and for simplicity we write simply “$G$ is an algebraic group” to mean that $G$ is an affine algebraic group in the sense of Milne’s book. Furthermore, apart from the exceptions already mentioned, we will only consider the case where the field $F$ has characteristic zero, so $G$ will be linear (meaning that it is isomorphic to a Zariski-closed $F$-subscheme of $\operatorname{GL}_{n,F}$ in a way that is compatible with the group operations on $G$ and $\operatorname{GL}_{n,F}$) and smooth, see [Mil17, Prop. 1.41, Th. 3.23, Cor. 4.10]. Because of the assumption of characteristic zero and that $G$ is affine, it is the same to consider these groups from the perspective of [Bor91] or [Spr98]. For any commutative $F$-algebra $R$, we write $G(R)$ for the set $\operatorname{Hom}(\operatorname{Spec}R,G)$ of $R$-points of $G$. The algebraic group $\operatorname{GL}_{1,R}$, i.e., the 1-dimensional torus with $R$-points $R^{\times}$, is typically denoted $\mathbb{G}_{\mathrm{m}}$. We sometimes write $\operatorname{Lie}(G)$ or $\mathfrak{g}$ for the Lie algebra of $G$, and similarly for other algebraic groups.
Mostly the symbol $T$ will denote a torus (most often a maximal torus) in a Lie or algebraic group $G$, and $W$ will denote the quotient $N_{G}(T)/T$.
We say that an algebraic group $G$ over a field of characteristic zero is reductive if its unipotent radical is trivial, i.e., if the identity component $G^{\circ}$ is reductive in the sense of [Mil17, §19c]. We say that $G$ is semisimple if it is connected and its radical (the maximal closed normal solvable subgroup of $G$) is the identity. Every semisimple group is reductive; a connected reductive group is semisimple if and only if its center is finite [Mil17, Prop. 19.10]. We say $G$ is simple ([Mil17, Def. 17.97] says “almost simple”) if it is semisimple and non-commutative, and every proper normal algebraic subgroup is finite.
A compact (real) Lie group $G$ can also be viewed as a reductive real algebraic group. The connected components of $G$ are the same, regardless of which viewpoint one takes. (The fact that the two notions coincide on the identity component is [Bor91, 24.6(ii)] and the more general statement follows, see [Con10].) Of course, for non-compact Lie groups this is typically false.
@@ 2. Serre’s bounds on the trace
Here $\operatorname{Ad}\colon K\to\operatorname{GL}(\operatorname{Lie}(K))$ is the adjoint representation, the action of $K$ on its own Lie algebra by conjugation. This result is proved in §14 as a consequence of the following more detailed result, which considers both the minima and maxima of $\operatorname{Tr}\operatorname{Ad}$ at the price of restricting to simple compact Lie groups. See [FG98] and [Fre01] for calculations of the value of $\operatorname{Tr}\operatorname{Ad}(x)$ for various $x$, in the case where $K$ is simple of type $E_{8}$.
@@
In addition to the results we prove here, Serre has recently published in [Ser23] his proofs of the other results announced in his 2004 note, so now proofs of all of the announced results are available.
@@ 3. Preliminary observations
Let $G$ be a simple compact Lie group with maximal torus $T$. There is a $\phi\in\operatorname{Aut}(G)$ that normalizes $T$ and acts as $\phi(t)=t^{-1}$ for $t\in T$ [Bou05, §IX.4.10, Prop. 17]. (It is even unique up to conjugation by an element of $T$.) Then $\phi$ interchanges the positive and negative roots, so the only nonzero diagonal entries of $\operatorname{Ad}(\phi)$ relative to a root space decomposition of $\operatorname{Lie}(G)$ come from its action on the torus and we have $\operatorname{Tr}\operatorname{Ad}(\phi)=-(\operatorname{rank}G)$. In the notation of Th. 2.3a, this proves that $\underline{m}(s)\leq-(\operatorname{rank}G)$.
@@
The element $g$ is contained in a maximal torus $T$ of $G$. Put $W$ for the Weyl group of $G$ with respect to $T$. Write $f$ as a composition $T\to T/W\to\mathbb{C}$. Suppose $g$ is regular. Then the Jacobian of the first map is invertible at $g$ by Steinberg’s Theorem as in [Ste60] or [Ste65, 6.1, 6.3]. Since $g$ is a critical point of $f$, it follows that the gradient of the map $T/W\to\mathbb{C}$ must vanish at the image of $g$, i.e., ii. ∎
We also need the following small elaboration of results of Steinberg. We provide a proof because we do not know a reference for it in the literature. Every algebraic group $G$ acts on itself by conjugation, and we write $\operatorname{Class}(G)$ for the (categorical) quotient of $G$ by this action (in order to avoid the notation $G/G$). As a scheme, for $G$ a reductive algebraic group over $F$, $\operatorname{Class}(G)$ is $\operatorname{Spec}F[G]^{G}$ and $F[G]^{G}$ is finitely generated.
@@
    If the torus $Z(G)^{\circ}$ is split and the derived subgroup $G^{\prime}$ of $G$ is simply connected, then
@@
To verify that the natural map in a is an isomorphism, it suffices to do so when $F$ is algebraically closed. If $G$ is semisimple, then this is a result of Steinberg, [Ste65, 6.4]. The proof is a reduction to that case.
Let $T^{\prime}$ be a maximal torus in the derived subgroup $G^{\prime}$ of $G$. The identity component $Z(G)^{\circ}$ of the center of $G$ is a torus and $T:=T^{\prime}\cdot Z(G)^{\circ}$ is a maximal torus of $G$. By conjugacy of maximal tori in $G$, we may assume that the given torus $T$ is this $T$. Set $H:=T^{\prime}\cap Z(G)^{\circ}$, $T^{+}:=T^{\prime}\times Z(G)^{\circ}$, and $G^{+}:=G^{\prime}\times Z(G)^{\circ}$. There is a natural diagram
@@
For the left vertical arrow, $T^{+}/W=T/W\times Z(G)^{\circ}$ and $\operatorname{Class}(G^{+})=\operatorname{Class}{G^{\prime}}\times Z(G)^{\circ}$, and these are isomorphic because $G^{\prime}$ is semisimple. That is, the left vertical arrow is an isomorphism. The top and bottom arrows are a quotient map by $H$, on which $W$ and $G^{+}$ both act trivially, so the right arrow is also an isomorphism, proving a.
@@
It thus remains to consider the action of $H$. Since $Z(G)^{\circ}$ is split, $\operatorname{Hom}(H,\mathbb{G}_{\mathrm{m}})$ is a finite abelian group with trivial action by the Galois group, and thus the action of $H$ induces an $F$-linear grading of $V_{G^{\prime}}$ by $\operatorname{Hom}(H,\mathbb{G}_{\mathrm{m}})$. In addition, the map $\operatorname{Hom}(Z(G)^{\circ},\mathbb{G}_{\mathrm{m}})\to\operatorname{Hom}(H,\mathbb{G}_{\mathrm{m}})$ is surjective, and thus we can rescale each homogeneous component by a character of $Z(G)^{\circ}$ to get an $H$-invariant affine factor. We thus get an isomorphism
$\operatorname{Class}(G)\cong(\mathbb{A}^{r^{\prime}}\times Z(G)^{\circ})/H\cong\mathbb{A}^{r^{\prime}}\times(Z(G)^{\circ}/H).$
Since $Z(G)^{\circ}$ is a torus, the quotient $Z(G)^{\circ}/H$ is also a torus (it is also the abelianization $G^{\text{ab}}$ of $G$), as required.
When $G^{\prime}$ is also split, $V_{G^{\prime}}$ is the span of fundamental characters, all of which are homogeneous for the grading by $\operatorname{Hom}(H,\mathbb{G}_{\mathrm{m}})$, and thus the procedure of the previous paragraph multiplies the fundamental characters by characters of $Z(G)^{\circ}$ to obtain new characters that descend to $G$. We thus get an isomorphism of the form
@@
The group $G=U(2)$ has simply connected derived subgroup $G^{\prime}=\operatorname{SU}(2)$, with $\operatorname{Class}(G^{\prime})\cong\mathbb{A}^{1}$. However, $Z(G)^{\circ}$ is a rank $1$ torus that is not split and we claim that $\operatorname{Class}(G)$ is not isomorphic to the product of an affine space and a torus.
@@ 4. Invariant derivations and critical points
The “only if” direction is trivial. So suppose $t$ is a critical point for $f|_{T}$. The identity component $G^{\circ}$ of $G$ is a union of conjugates of $T$, and thus in particular the conjugates of $\mathfrak{t}$ span $\mathfrak{g}$. For $x\in\mathfrak{g}$, there exists $g_{1},\ldots,g_{n}\in G$ and $y_{1},\ldots,y_{n}\in\mathfrak{t}$ such that $\sum_{i}(\operatorname{Ad}g_{i})y_{i}=x$ and we have
$xf(t)=\sum_{i}((\operatorname{Ad}g_{i})y_{i})f(t)=\sum_{i}y_{i}(g_{i}^{-1}f)(t)=\sum_{i}y_{i}f(g_{i}tg_{i}^{-1}).$
By hypothesis, the summand in the last term equals $y_{i}f(t)$, which is zero because $t$ is a critical point for $f|_{T}$. ∎
@@
$x$ is a critical point for $f$ if and only if $(Df)(x)=0$ for every derivation $D$ (not necessarily $H$-invariant), so the “only if” direction is trivial and we assume $(Df)(x)=0$ for every $H$-invariant $D$. Let $D_{0}$ be any derivation. Since $H$ is finite, there is a $\pi\in F[X]$ such that $\pi(x)=1$ and $\pi$ vanishes on the other elements of the $H$-orbit of $x$. Put $H_{x}$ for the centralizer of $x$ in $H$ and define
@@ 5. Derivations on $T/W$ versus $W$-invariant derivations on $T$
In this section, we work in the context of a connected reductive algebraic group $G$ over $\mathbb{C}$ with a maximal torus $T$ and Weyl group $W:=N_{G}(T)/T$. We relate the $W$-invariant derivations on $T$ — which we denote by $(\operatorname{Der}\mathbb{C}[T])^{W}$ — to derivations on $T/W$, i.e., $\operatorname{Der}\mathbb{C}[T/W]$.
Any $W$-invariant derivation on $T$ takes $W$-invariant functions to $W$-invariant functions, so induces a derivation on $T/W$, giving us a natural map $(\operatorname{Der}\mathbb{C}[T])^{W}\to\operatorname{Der}\mathbb{C}[T/W]$. Every derivation on $\mathbb{C}[T/W]$ extends uniquely to its function field $\mathbb{C}(T/W)$, and hence to that field’s finite separable extension $\mathbb{C}(T)$. In particular, a derivation on $T$ is determined by its values on $\mathbb{C}[T/W]$, so this map is injective. This lets us identify $(\operatorname{Der}\mathbb{C}[T])^{W}$ with a submodule of $\operatorname{Der}\mathbb{C}[T/W]$.
@@
$(\operatorname{Der}\mathbb{C}[T/W])/(\operatorname{Der}\mathbb{C}[T])^{W}\cong\mathbb{C}[t]/(t^{2}-4)$
and in particular that $(\operatorname{Der}\mathbb{C}[T])^{W}$ is a proper submodule of $\operatorname{Der}\mathbb{C}[T/W]$.
It turns out that when $T/W$ is sufficiently nice (e.g., when the derived subgroup $G^{\prime}$ is simply connected), not only can we show that the relevant submodule of $\operatorname{Der}\mathbb{C}[T/W]$ is a free submodule, but the proof gives us a reasonably natural way to construct a basis. (A similar construction appears in §B.2 of [OT13], although that material concerns the linear action of $W$, the quotient of a euclidean space by the reflection representation of $W$.)
@@
We take a moment to review some background on differential operators on an affine variety $X$ over $\mathbb{C}$.
@@
A differential operator of order 0 is $\mu_{f}\colon\mathbb{C}[X]\to\mathbb{C}[X]$ given by multiplication by an element $f\in\mathbb{C}[X]$. A differential operator $\mathcal{D}$ of order $k+1$ is a $\mathbb{C}$-linear map $\mathbb{C}[X]\to\mathbb{C}[X]$ such that, for each $g\in\mathbb{C}[X]$, the map $f\mapsto\mathcal{D}(gf)-g\mathcal{D}(f)$ is a differential operator of order $k$.
The collection of differential operators of order $k$ is denoted $\operatorname{Diff}^{k}_{\mathbb{C}[X]/\mathbb{C}}(\mathbb{C}[X],\mathbb{C}[X])$ in [Sta18, Tag 09CH], and it is a $\mathbb{C}$-module. A differential operator of order $k$ is a differential operator of order $K$ for all $K\geq k$. A differential operator of order 1 is a derivation if and only if it annihilates 1 [Sta18, Tag 09CM].
@@
Let $\mathcal{D}$ be a second-order differential operator on an affine variety $X$ over $\mathbb{C}$. Then for $f,g\in\mathbb{C}[X]$:
@@
    $[\mu_{f},\mathcal{D}]+\mu_{\mathcal{D}(f)}-\mu_{\mathcal{D}(1)f}$ is a derivation on $X$.
@@
    The pairing $(f,g)\mapsto\mathcal{D}(fg)-f\mathcal{D}(g)-\mathcal{D}(f)g+fg\mathcal{D}(1)$ is symmetric and is a derivation on $X$ in each slot. It is identically zero if $\mathcal{D}$ is a first-order differential operator.
@@
For b, the pairing is evidently symmetric and $\mathbb{C}$-linear in each slot. If $\mathcal{D}$ is first order, then there is an $f^{\prime}\in\mathbb{C}[X]$ such that $\mathcal{D}(fg)-f\mathcal{D}(g)=f^{\prime}g$ for all $g$. Then $\mathcal{D}(f)g-fg\mathcal{D}(1)=f^{\prime}g$ as well and the pairing is zero. Finally, the the image of $(f,g)$ is $-1$ times the derivation in a, so it is a derivation in $g$. ∎
@@
We now return to the situation of $G$, $T$, and $W$ from the beginning of this section. Let $A$ be a nonsingular $W$-invariant quadratic form on $\mathfrak{t}^{*}$. We may view this as an element of $(S^{2}\mathfrak{t})^{W}$, i.e., a Weyl-invariant element of the symmetric square. Since elements of $\mathfrak{t}$ are naturally translation-invariant derivations, $A$ corresponds naturally to a second-order differential operator $\mathcal{D}_{A}$ on $\mathbb{C}[T]$, invariant under $W$ and under translation by $T$.
@@
$\Omega_{\mathbb{C}[T/W]}\to\operatorname{Der}\mathbb{C}[T/W]$
induced by $M_{A}$ is $(\operatorname{Der}\mathbb{C}[T])^{W}$.
@@
$\Omega_{\mathbb{C}[T]}\cong\operatorname{Der}\mathbb{C}[T],$
@@
$\Omega_{\mathbb{C}[T]}^{W}\cong(\operatorname{Der}\mathbb{C}[T])^{W}.$
@@
$\Omega_{\mathbb{C}[T/W]}\to\operatorname{Der}\mathbb{C}[T/W]$
@@
$\Omega_{\mathbb{C}[T/W]}\to\Omega_{\mathbb{C}[T]}^{W}\cong(\operatorname{Der}\mathbb{C}[T])^{W}\hookrightarrow\operatorname{Der}\mathbb{C}[T/W]$
in which the maps on either side are natural. Moreover, by [Kni73, Cor. to Th. II.3], the first natural map is an isomorphism, and thus the image of the map induced by $M_{A}$ is $(\operatorname{Der}\mathbb{C}[T])^{W}$ as required. ∎
@@
If the derived subgroup $G^{\prime}$ of $G$ is simply connected, then $(\operatorname{Der}\mathbb{C}[T])^{W}$ is a free module over $\mathbb{C}[T]^{W}$.